*Present:* Bucko, Bruce, Mullet, Si, Polar, Dan, Fluff_eei and Sadiq
* Previous Minutes (BF Planning Meeting)
* Money matters
* Gaming Events
* New Gaming Survey
* Communication with Tech Team and Members
* Server Upgrade Status
* Molotov Status
* Flyers and Promotional Booklet
* LAN Planning
* Clothing Orders
* Actions Points Previous Meetings
* Date for Next Meeting
h4. Previous Minutes
Accepted by all.
h4. Money Matters
Bruce stated that no one knew the budget for next year (as of 23rd April).
h4. Communication with Tech Team and Members
Mullet stated that whenever upgrades or installations happened the Tech Team needed to inform members of the upgrade date/time and when the servers may be down. This had not happened for the recent kernel upgrade. This could easily be fixed by an email to TechTeam or to members in the case of a serious update being applied. Whether this happens late at night or early in the day an email is an important way to communicate the changes to members.
*Action* Tim is to ensure TechTeam *always* inform members of upgrades/software changes or planned downtime.
h4. Molotov Status
Mullet can confirm that the server is now working. The kernel was rolled back to the previous (pre-Etch) version to prevent the /var errors that have been occurring.
*Action* TechTeam/Tim to edit GRUB configuration to ensure it boots from working kernel - not Etch, until problems and solution is diagnosed.
h4. New Servers Installation Status
Mullet can confirm Backus is up and running but that Tim has not migrated any software (as yet).
Si asked which sub-TechTeam teams had responsibility for this - apparently most sub-teams had completed their role with only the Wiki team left to install software on Backus.
Sadiq asked if Bong was still down, apparently it is. Both Mullet and Polar contributed that various components and Bong's IP address were being used by Backus.
Mullet stated that Backus was intended to replace Twilight and Bermuda for now.
Sadiq said that the non-Naur server was to be the vserver testing ground until the summer when the upgrade will be completed.
*Action* Tim and TechTeam to report to Exec on how upgrade is going and what parts are completed (as next Exec meeting).
*Action* Si is to talk to Bruce post exams about selling RAID cards.
Sadiq asked whether Tim had spoken to Chris Tilbury. Mullet stated that Tim has attended a meeting and said the Society would forward a list of ports to ITS as soon as possible but that the SMTP/E-Mail server port would not be opened. Bucko reminded the Exec that there is a table in MySQL which he setup which contains the ports used by the Society for gaming.
*Action* Tim to email Chris Tilbury ASAP.
h4. Gaming Survey
Mullet stated that Skumby had completed a gaming survey in previous years (results at skumby.uwcs.co.uk). We need to complete a survey this year as gaming was restarting.
*Action* Fluff_eei to produce a survey for members.
*Action* Mullet and Fluff_eei to find games in cupboard (after meeting)
h4. Gaming Events
Following Si's email to Roland Wilson, Mullet and Si had a meeting with the CS-SSLC to obtain permission for gaming events to be held in the department. They reported that the meeting seemed to go well and that the SSLC had seemed positive provided that the Society worked around deadlines (when resources may become heavily used).
Si said that until we actually have a DCS gaming username the gaming cannot resume. Mullet stated that he wanted to get this quickly if possible.
Dan said that after conversations with some ITS people, there was the impression amongst some ITS staff that CompSoc had not requested the use of the re-fitted Computer Rooms.
*Action* Andy (to talk to Bucko) and send email to ITS staff regarding new rooms.
*Action* Mullet and Sadiq to find out about Power.
Polar has added LANs to website but the information needs updating.
Mullet told the meeting that he had talked to Stig at Bloomberg who seemed interesting in sponsoring the Programming Competition.
*Action* Mullet to email Stig.
Polar said IBM had not replied.
h4. Flyers and Promotional Materials
Sadiq stated that in previous years the Society had send out flyers with the start of year Freshers ITS packs. We needed to do this again.
*Action* Andy to contact Bea Mistry at ITS with regards this.
*Action* Si to produce a Poster for gaming.
Sadiq raised the issue of some new email groups to spread requests that new members might have. For instance, exec, academic, social, gaming and members.
*Action* Tim to change website so by default members subscribe to all of the above and then have to unsignup.
*Action* Polar to remove poster in SU North Building.
*Action* Polar to put up new poster with URLs for CompSoc
h4. LAN Planning
Fluff_eei said that Loughbrough apparently think they are invited to the LAN.
Si said following conversations with Steph Wongsam and SU, they will not be covered by *any* insurance.
Bruce said that Loughbrough would probably bring around 15 people.
*Action* Fluff_eei to contact Loughbrough CS Society - are they still coming?
*Action* Mullet and Sadiq to look at Power issues for LANs.
Sadiq told the exec that ADM who used to produce t-shirts for the society currently own the digitised version of the logo and they would not let us have it. So he has asked his supplier to look into digitising the logo again and what prices he might charge.
*Action* Sadiq to bring price lists to exec.
h4. Crypt and Graveyard
Sadiq will pick them up from Pete.
h4. Action Points
*Action* Si to tell Mullet about BBQ Week 3 event
*Action* Fluff_eei to email members regarding a film in week 10.
*Action* Sadiq to send email asking who wants to go to Bingo week 10.
*Action* Polar to talk to icStatic regards Gaming
*Action* Polar and Si to do Guidebook
*Action* Si to type up Bucko's papers relating to Gaming
*Action* Sadiq to install Process Accounting
*Action* Si to email membership with details of failures
*Action* TechTeam are to get backups working as soon as possible since upgrades may require backed-up data at some point
*Action* Each person setting up a new service on Backus/Not-Naur must create documentation of what they have done.
*Action* Bucko to poke Fred regards Planet.
*Action* Bruce to look into how we pay for Cedega.
h4. Next Meeting
To be held Tuesday, 15:00.